OpenStudy (anonymous):

If x^2+y^2 =25 and dy/dt =6, find dx/dt when y=4

OpenStudy (anonymous):

2x dx/dt +2y dy/dt =0 x dx/dt +y (6)=0 dx/dt= -6y/x dx/dt= -6*4 /x dx/dt=-24/x also from the given equation x^2= 25-16=9 or x=3 using this we get dx/dt=-24/3 =-8

OpenStudy (anonymous):

great answer, thanks for breaking it down for me

OpenStudy (anonymous):

no worries